Determine the form of a particular solution for each differential equation. No need to solve for the general solution of the differential equation.


1. y'' − 3y' + 2y = xe^x + 1


1
Expert's answer
2022-02-22T23:19:05-0500

Corresponding homogeneous differential equation


"y''-3y'+2y=0"

Characteristic (auxiliary) equation


"r^2-3r+2=0"

"r_1=1, r_2=2"

Find a particular solution in form "y_p=x(Ax+B)e^x+C."


"y_p=Ax^2e^x+Bxe^x+C"


"y_p'=Ax^2e^x+2Axe^x+Bxe^x+Be^x"


"y_p''=Ax^2e^x+4Axe^x+2Ae^x+Bxe^x+2Be^x"

"Ax^2e^x+4Axe^x+2Ae^x+Bxe^x+2Be^x"

"-3Ax^2e^x-6Axe^x-3Bxe^x-3Be^x"

"+2Ax^2e^x+2Bxe^x+2C=xe^x+1"


"-2A=1"

"2A-B=0"

"2C=1"

"A=-1\/2, B=-1, C=1\/2"

The particular solution is


"y_p=-\dfrac{1}{2}x^2e^x-xe^x+\dfrac{1}{2}"